OpenStudy (buttermequeasy):

How many solutions does the equation 2|x − 3| + 4 = 4 have? Zero One Two Infinitely Many

OpenStudy (buttermequeasy):

I'm pretty sure it's two.

OpenStudy (buttermequeasy):

???

OpenStudy (precal):

If you solve it, the solution is 3. Sub x=3 \[2\left| 3-3 \right|+4=4\] \[2\left| 0 \right|+4=4\] 0+4=4 4=4

OpenStudy (buttermequeasy):

Yes, I know, but how many solutions does this one equation have? More than one or just one?

OpenStudy (buttermequeasy):

???

OpenStudy (anonymous):

2! Bc when u solve for an absolute value problem, you set up two equations

OpenStudy (anonymous):

ONE! 2|x − 3| + 4 = 4 given 2|x − 3| = 0 subtr 4 |x − 3| = 0 divide by 2 x=3 is the only solution

OpenStudy (anonymous):